OpenStudy (anonymous):

pi.i was helping me with a problem. I had to add to it so need more help please.

OpenStudy (anonymous):

sure what do you need help with

OpenStudy (anonymous):

i added it to my last post. I can try and copy/paste it

OpenStudy (anonymous):

Ok this is what I have h=(4,38) &p=(13,95) I need to fine the point slope of an equation of the line through the 2 ordered pairs. My options are: a.) p=19/3h+38/3 b.) p-4=3/19(h-38) c.) p-38=19/3(h-4) d.) p-38=-19/3(h-4) so which one is the answer?

OpenStudy (anonymous):

so now it is asking to use that formula to estimate the production of an 18 foot tree anda 20 foot tree

OpenStudy (anonymous):

the formula is p-38=19/3(h-4)

OpenStudy (anonymous):

hey, do you still need help?

OpenStudy (anonymous):

ya sure do

OpenStudy (anonymous):

your question is asking for point-slope form of the equation?

OpenStudy (anonymous):

no it is asking to use the formula p-38=19/3(h-4) to estimate the production of an 18 foot and 20 foot tree.

OpenStudy (anonymous):

plug each height in seperately in the formula for h and solve for p